# FloorplanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations

### v2

@@ 3.5 Evaluation Protocol and Scoring
Each question in FloorplanQA is paired with a reference answer computed directly from the symbolic layout, enabling fully automated and deterministic evaluation of model outputs. Depending on the response type, correctness is assessed using numeric comparison with fixed tolerances, string matching, or geometric validation checks.
For numerical questions (e.g., distances, areas, angles), predictions are accepted if they fall within a relative error threshold. Sequence outputs are evaluated for both format and semantic correctness. Rather than requiring exact coordinate matches, we validate paths using geometric plausibility conditions of collision avoidance and sufficient proximity to a ground-truth trajectory. Deviation thresholds are set conservatively to tolerate minor geometric variations without crediting qualitatively wrong routes.
To differentiate reasoning failures from extraction or formatting issues, we apply a regex-based parsing pipeline, covering a fixed set of expected answer patterns (e.g., ‘*Final answer*’ tokens in lower case or with surrounding symbols). If no answer is produced, or if the extracted content does not match a valid format, we count the response as an error (which is also considered incorrect). We provide a detailed breakdown on accuracy in Appendix E. In our evaluations, the proportion of invalidly formatted answers is below 1%.
We also explicitly track cases where no answer is returned due to truncation (API: stop_reason = Token Limit), a failure mode that disproportionately affects reasoning-heavy models. To account for this, we report in Appendix I both the percentage of responses truncated by token limits and an adjusted accuracy computed only over valid (non-truncated) answers. These adjusted accuracies can be interpreted as an approximate upper bound on overall dataset performance, since with unlimited token budgets models could in principle reach that quality.
@@ 3.6 Model Inference Setup
All models are queried using standard chat-based completion APIs, with prompts constructed as described above. We evaluate both large and mid-size models, including reasoning and standard variants. Large reasoning-oriented and standard models are allocated up to 12,288 tokens per completion, enabling them to process long layout descriptions and produce multi-step outputs. Mid-size models, including the GPT family variants optimized for speed and Qwen3-30B, are limited to 8,192 tokens. These budgets reflect our observations that larger models generate longer intermediate justifications and thus consume more tokens.
GPT-5 is evaluated under a distinct configuration: its reasoning intensity can’t be disabled, so we run it with reasoning and verbosity set to “low,” with a maximum output length of 4,096 tokens, while GPT-5-mini is run with reasoning and verbosity set to “medium,” with a maximum output length of 8,192 tokens.
No model receives fine-tuning or prompt adaptation specific to FloorplanQA. All prompts are zero-shot, with the system message and output formatting constraints held fixed. For each model, we use the default inference configuration provided by its vendor, with temperature set to 0. The only exception is GPT-5, for which the temperature cannot be modified and defaults to 1.
@@ 4 Results
We evaluated the performance of the model on a dataset of 2,000 layouts, consisting of 600 kitchens, 600 living rooms, 600 bedrooms, and 200 additional layouts from HSSD. Each layout is paired with one question sampled from a pool of 8 parameterized templates, filtered by room applicability. This yields 16,000 layout–question pairs. The full taxonomy is shown in Table 2.
We evaluate fifteen models, spanning a wide range of parameter scales, architectures, and training regimes. The reasoning-oriented models include GPT-5 (OpenAI, 2025a), GPT-OSS-120B (OpenAI, 2025b), DeepSeek-R1-0528 (DeepSeek-AI, 2025), GPT-5-mini, Gemini Flash 2.5 (Google DeepMind, 2025), GPT-OSS-20B, and Qwen3-30B-A3B-Thinking-2507 (Yang et al., 2025). The general-purpose (standard) models include Claude Sonnet 4 (Anthropic, 2025), GPT-4.1 (OpenAI et al., 2024), Moonshot Kimi-K2-Instruct (Kimi Team, 2025), Qwen3-Coder-480B-A35B-Instruct, Qwen3-235B-A22B-Instruct-2507, GPT-4.1-mini, Qwen3-30B-A3B-Instruct-2507, and Devstral-Small-2505 (MistralAI, 2025). All models are evaluated in identical zero-shot conditions using standardized prompts and serialized layout inputs, as described in Section 3. Complete results disaggregated by question type, room type, and model are provided in Appendix E.
@@ 4.1 Quantitative Results
We begin by aggregating accuracy across models and question types for both reasoning and general model families. Figure 2 summarizes accuracy: the top row shows general models, and the bottom row shows reasoning models. In each figure, the left panel summarizes accuracy by model, and the right panel summarizes accuracy by question across room types.
Kitchens lead across models because overlaps are rare, so most queries are straightforward. Scores on HSSD tend to lag behind the other room types; irregular, non-axis-aligned geometry and denser overlap make these layouts more demanding, see Appendix D for detailed statistics. Bedrooms and living rooms are mid-tier and nearly equal, lying between kitchens and HSSD.
Comparing model families, general models struggle when many overlaps must be merged; they often treat object areas independently (no union), which hurts Free Space and Max Box and carries over into path planning. Reasoning models handle those cases with unions and rotations better, so they show gains on Free Space and Max Box. A practical limitation remains: Gemini Flash 2.5 and DeepSeek-R1 hit token limits on larger layouts, which drops scores, especially in HSSD.
Task difficulty follows a consistent pattern in terms of accuracy. Metric-category questions, such as Pair Distance and View Angle, achieve the highest accuracies. Repositioning, Visibility, and Placement yield mid-range accuracies. Max Box and Free Space benefit most from reasoning-oriented models, yet their accuracies remain low and are comparable in difficulty to Shortest Path. Accuracy decreases as object count and overlap density increase, most notably on HSSD layouts. Detailed analysis on each question category is provided in Appendix G
@@
Figure 2: Accuracy of general (top) and reasoning (bottom) models. Left: by model. Right: by question. Each column corresponds to a specific room type represented in our dataset: Kitchens, Living Rooms, Bedrooms (synthetic subsets), plus HSSD layouts.
4.2 Ablation
To evaluate the robustness of layout interpretation under alternate encodings, we perform a format ablation that replaces the standard JSON layout with a semantically equivalent XML version. This substitution preserves all geometric and object-level content while modifying only the syntax and structural serialization. The prompt level and semantic ablations are described in Appendix, Section F. To avoid recomputing the full suite, we focus on the most variance-sensitive question types for HSSD layouts, as shown in Appendix F (Figs. 5 and 6): View Angle, Visibility, and one task from a different category—Repositioning (Dynamic). We apply each input-format ablation to three representative models: two reasoning models (GPT-OSS-120B, GPT-OSS-20B) and one large general model (Qwen3-235B-A22B) for the same subset of HSSD layouts. As shown in Table 3, accuracy is largely stable in both formats for most categories; the GPT-OSS models change minimally, while Qwen shows modest fluctuations. This suggests that these models encode layout semantics in a manner that is relatively invariant to low-level representation details.
Model Repositioning View Angle Visibility
GPT-OSS-120B 60.5 → 59.0 74.0 → 74.0 70.0 → 72.0
GPT-OSS-20B 40.0 → 39.0 37.5 → 38.5 45.5 → 43.5
Qwen3-235B-A22B 39.0 → 42.0 50.5 → 54.0 70.5 → 65.5
@@ 5 Conclusion
We introduced FloorplanQA, a benchmark for spatial reasoning over symbolic 2D layouts aligned with architecture and robotics practice. We evaluated 15 language models (8 general, 7 reasoning) on 2,000 layouts (1,800 synthetic; 200 semi–real HSSD) across tasks ranging from metric queries to visibility, placement, and shortest path with clearance.
Empirically, metric and simple visibility queries are reliable; kitchens score highest because overlaps are rare. HSSD layouts are more demanding—irregular, non–axis-aligned shapes and dense overlaps expose weaknesses such as centroid miscalculation and missed unions. Reasoning models improve notably on Free Space and Max Box by handling overlaps and rotations more consistently, while general models often subtract object areas independently and fail under heavy overlap. Shortest Path is sufficiently challenging because it requires multiple correct steps (clearance buffering, collision checks, and path search), where errors compound.
These results indicate that current LLMs lack sufficiently robust internal geometric representations for complex spatial inference. Two complementary directions follow. Near term: hybridize with external geometric solvers or symbolic planning modules—set operations (unions/differences), centroid via shoelace, clearance buffering, oriented rectangle search, and A* path planning—to compensate for the models’ weaknesses in collision avoidance and clearance reasoning. Longer term: train with explicit spatial constraints and harder distributions (irregular, overlap–heavy layouts), and include constraint–violation exemplars and geometry-aware objectives so models learn to maintain coherence under rotation, clearance, and union operations in design-oriented tasks.
@@ Acknowledgments
The work is supported by funding from King Abdullah University of Science and Technology (KAUST)—Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program
@@ Appendix A Technical Appendix and Supplementary Material for FloorplanQA
This appendix details the dataset generation pipeline, dataset statistics, and the full evaluation of FloorplanQA across reasoning and general models. It also includes ablation studies, analysis of failure cases, and the exact prompts used for data generation and question asking.
@@ B.1 Room Shape Generation
Room Type
Shape Distribution
(Rect / L-shaped / Open)
Size Categories (in m2)
(Small / Medium / Large)
@@ Appendix C Details on HSSD layouts selection
During 2D projection, we correct mislabeled windows and doors, snap nearly collinear edges, and resolve self-intersections. To smooth rectilinear artifacts and irregular edges in box-like footprints, we apply alpha-shape–based convex hulls, then recompute centroids and areas on the cleaned geometry.
@@ Appendix D Room Statistics
Table 5 reports summary statistics for the layouts. Kitchens are typically the smallest spaces, with relatively few objects and overlaps, but a high density due to their compact geometry. Living rooms are the largest, with slightly more objects overall but lower density, reflecting their open layout. Bedrooms fall in between, with similar object counts to kitchens but more frequent overlaps.
@@
Avg. # of Overlaps 0.52 1.82 1.52 4.39
@@ E.1 Detailed Accuracy by full dataset
Tables 6 and 7 report the full per-model accuracy matrix for the base and reasoning model groups, respectively. Each table covers nine question categories across four room types, resulting in a total of 36 rows. The path task is assessed using two complementary metrics: validity and Fréchet Distance. Together, these question types span a range of challenges, covering both reasoning-heavy and functionally grounded tasks. The base group includes eight general-purpose models, while the reasoning group includes seven models with explicit reasoning capabilities. For every model, question type accuracy is computed over a fixed set of 600 synthetic layouts (Kitchens, Living Rooms, Bedrooms) and 200 HSSD layouts, ensuring that results are directly comparable across models.
E.2 Token-Limit Analysis and Valid-Only Accuracy
First, Tables 10 and 12 quantify the fraction of responses that were terminated due to the TOKEN LIMIT stop reason. This failure mode is particularly relevant for reasoning-oriented models, which often generate longer chain-of-thought outputs.
Second, Tables 11 and 13 report accuracy over completed answers only, excluding truncated outputs (e.g., token-limit terminations). This metric isolates models’ reasoning performance on successfully produced, well-formed responses.
@@ E.2 Token-Limit Analysis and Valid-Only Accuracy
Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small

### v3

@@ 3.4 Evaluation Protocol and Scoring
Each question in FloorplanQA is paired with a reference answer computed directly from the symbolic layout, enabling fully automated and deterministic evaluation of model outputs.
Answer extraction. To differentiate reasoning failures from formatting issues, we apply a regex-based parsing pipeline covering expected answer patterns (e.g., ‘*Final answer*’ tokens). If no answer is produced or the extracted content does not match a valid format, we count the response as an error. The proportion of invalidly formatted answers is below 1%. Per-model truncation rates and invalid-format proportions are summarized in Table 7 (Appendix E.3).
Truncation tracking. We track cases where no answer is returned due to token limits (API: stop_reason = Token Limit), a failure mode that disproportionately affects reasoning-heavy models. We report both the raw, detailed accuracy breakdown by model, room type, and question type (Appendix E, Tables 5, 6) and the adjusted accuracy computed only over non-truncated responses (Appendix K). The adjusted scores serve as an upper bound on performance, since larger token budgets could enable more complete solutions. However, truncation occurs primarily on the most challenging questions, so valid-only accuracy may inflate performance and should not be interpreted as a standalone metric.
@@ 3.5 Model Inference Setup
We evaluate both large and mid-size models, including reasoning and standard variants. Large reasoning-oriented and standard models are allocated up to 12,288 tokens per completion, enabling them to process long layout descriptions and produce multi-step outputs. Mid-size models, including the GPT family variants optimized for speed and Qwen3-30B, are limited to 8,192 tokens. These budgets reflect our observations that larger models generate longer intermediate justifications and thus consume more tokens. GPT-5 is evaluated under a distinct configuration, since its reasoning intensity cannot be disabled. We run GPT-5 with reasoning enabled and “low” verbosity and limit its output budget to 4,096 tokens, counting both reasoning and generated tokens, while accounting for its higher computational cost. GPT-5-mini is evaluated with “medium” reasoning and verbosity and an 8,192-token output budget.
For each model, we use the default inference configuration provided by its vendor, with temperature set to 0. The only exception is GPT-5 family, for which the temperature parameter is not user-configurable and defaults to 1 when reasoning is enabled.
@@ 4 Results
We evaluate fifteen models on the full benchmark: 2,000 layouts (600 kitchens, 600 living rooms, 600 bedrooms, 200 HSSD) with eight questions each, yielding 16,000 layout–question pairs. The reasoning-oriented models include GPT-5 (OpenAI, 2025), gpt-oss-120B (Agarwal et al., 2025), DeepSeek-R1-0528 (DeepSeek-AI, 2025), GPT-5-mini, Gemini Flash 2.5 (Google DeepMind, 2025), gpt-oss-20B, and Qwen3-30B-A3B-Thinking-2507 (Yang et al., 2025). The general-purpose models include Claude Sonnet 4 (Anthropic, 2025), GPT-4.1 (OpenAI et al., 2023), Moonshot Kimi-K2-Instruct (Kimi Team, 2025), Qwen3-Coder-480B-A35B-Instruct, Qwen3-235B-A22B-Instruct-2507, GPT-4.1-mini, Qwen3-30B-A3B-Instruct-2507, and Devstral-Small-2505 (Mistral AI and All Hands AI, 2025). All models are evaluated in identical zero-shot conditions. The full taxonomy is shown in Table 1; complete per-model results are provided in Appendix E.
@@ 4.1 Quantitative Results
Room type difficulty. Figure 3 summarizes accuracy across models and question types, with general models in the top row and reasoning models in the bottom row. Kitchens yield the highest accuracy due to simpler geometry and fewer object overlaps: on average, each kitchen layout contains 0.52 overlapping object pairs (excluding rugs), compared to 1.52–1.82 overlapping pairs in bedrooms and living rooms (Appendix D, Table 4). HSSD layouts are more complex due to non-axis-aligned polygons and substantially more overlap (4.39 overlapping pairs per layout, excluding rugs), which increases the difficulty of union, clearance, and collision reasoning. Bedrooms and living rooms fall in between. Model rankings (left panel) mainly remain consistent across synthetic and HSSD layouts, suggesting that synthetic layouts test the same spatial reasoning skills as human-authored scenes.
@@
Figure 3: Accuracy of general (top) and reasoning (bottom) models. The left panel summarizes accuracy by model, averaged across all question types. The right panel summarizes accuracy by question, averaged across all models within the respective general or reasoning family. Each column corresponds to a specific subset in our dataset: HSSD layouts, plus our Kitchens, Living Rooms, and Bedrooms.
Task difficulty hierarchy. Three distinct tiers emerge from analysing the right panel of Figure 3. Metric tasks (Pair Distance, View Angle) achieve 75–95% averaged accuracy on synthetic layouts, where objects are axis-aligned boxes. On HSSD layouts, accuracy drops to 35–60%: models frequently fail to compute centroids of complex polygons, making computational errors when applying the shoelace formula (see Appendix I.1, I.4), compared to simple averaging for 4-point boxes. Constraint reasoning tasks (Placement, Visibility, Repositioning) yield 60–85% averaged accuracy; performance degrades when multiple geometric constraints must be jointly verified. For Placement and Visibility, baseline-guessing is ruled out by balanced ground truth (49.9% True) and low empty-list frequency (¡20%), respectively (see Appendix I.5, I.7, I.2). Optimization and planning tasks (Free Space, Max Box, Shortest Path) achieve only 5–45% averaged accuracy—tasks requiring geometric unions, search, or multi-step collision-free reasoning remain largely unsolved. One exception is Free Space on Kitchens, yielding over 90% accuracy due to minimal object overlaps; in contrast, other layouts require computing unions of overlapping object pairs (e.g., TV and TV stand, nightstand and lamp) to correctly calculate non-occupied area (see Appendix I.3, I.8, I.6).
Reasoning vs. general models. Reasoning-focused models show notable improvements on geometric union tasks, gaining +10–40% on Free Space and Max Box compared to general models on generated layouts. This suggests that extended inference computation helps with complex geometric operations. However, gains on path planning remain modest. A practical limitation affects some reasoning models: Gemini Flash 2.5 and DeepSeek-R1 frequently hit token limits on larger layouts (30–50% truncation; see Table 7), reducing their reported scores. Valid-only accuracy (Appendix K, Tables 13 and 15) shows these models perform competitively when responses are completed successfully.
4.2 Ablations
Input format. To test robustness to layout encoding, we replaced JSON with semantically equivalent XML for a subset of HSSD layouts across three tasks (View Angle, Visibility, Repositioning) and three models (gpt-oss-120B, gpt-oss-20B, Qwen3-235B). Accuracy remained stable across formats ($\pm$3 percentage points; see Appendix H.1, Table 9), suggesting models encode layout semantics independently of serialization syntax.
Prompt and semantic variation. We tested robustness to object reference substitution (different object names, same template) and semantic perturbation (permuted object labels, identical geometry). Object reference substitution shows minimal effect on accuracy (Appendix H.2). Semantic perturbation reveals a task-dependent pattern: View Angle and Visibility remain stable when object labels are swapped, but Repositioning accuracy drops sharply (e.g., 60.5% → 40.0% for gpt-oss-120B). We hypothesize that for movement-related tasks, models rely on linguistic associations rather than geometric properties—for instance, when asked to move a ”sofa” whose label has been swapped with ”chair,” models sometimes search for furniture that semantically resembles a sofa rather than using the object’s actual geometry (Appendix H.3).
Tool augmentation. To isolate whether failures stem from arithmetic errors or deeper reasoning limitations, we evaluated GPT-4.1 with an integrated Python Code Interpreter. Table 2 shows results on HSSD layouts. Additional results for generated layouts and GPT-4.1-mini model can be found in Appendix F, Table 8.
Table 2: Accuracy on HSSD layouts with and without Python Code Interpreter (GPT-4.1).
Task Raw Tools $\Delta$
Pair Distance 56.0 99.0 +43.0
View Angle 55.0 96.0 +41.0
Visibility 46.5 86.5 +40.0
Repositioning 47.0 83.5 +36.5
Placement 64.5 95.0 +30.5
Free Space 16.0 44.0 +28.0
Max Box 7.5 3.0 -4.5
Shortest Path 25.0 12.5 -12.5
Tool augmentation substantially improves arithmetic-heavy tasks: Pair Distance (+43 points), View Angle (+41 points), Visibility (+40 points). However, performance on optimization and planning tasks decreases: Max Box (-4.5 points) and Shortest Path (-12.5 points). Inspection of failure cases (Figure 8) reveals that tool-augmented models generate Python code with incorrect algorithms—treating boundary contact as collision, or performing incomplete search over rectangle orientations. This indicates that the bottleneck for complex spatial tasks is algorithmic reasoning about spatial constraints, not numerical precision.
Vision-language input. We also evaluated VLM inputs using rendered floorplan images alongside symbolic layouts. Results (Appendix F, Table 8) show selective improvements on synthetic layouts but no consistent advantage over symbolic input alone, suggesting structured representations already provide a strong baseline for these tasks.
@@ 5 Conclusion
We introduced FloorplanQA, a benchmark for evaluating spatial reasoning in LLMs over symbolic 2D layouts representative of architectural and planning contexts. We evaluated 15 language models (8 general, 7 reasoning) on 2,000 layouts (1,800 synthetic; 200 semi–real HSSD) across 8 types of questions spanning metric queries, visibility, placement, and collision-free path estimation.
Key findings. Our evaluation (Section 4) reveals a consistent capability gap: while models reliably compute some geometric properties (distances, angles), they struggle with constraint satisfaction, geometric unions, and multi-step spatial planning. Three systematic failure modes emerge: (1) Models frequently behave as if object areas were independent rather than unioned (that is, they compute the sum of areas rather than the union), which aligns with low accuracy on Free Space and Max Box (Appendix D, Fig. 17). (2) Performance drops on HSSD layouts with non-axis-aligned, high-vertex polygons, and qualitative analysis indicates that many Distance/Angle errors originate in centroid computation (Appendix I.1, I.4). (3) Shortest Path remains difficult, especially in cluttered scenes where buffering and obstacle interactions narrow corridors (Appendix I.6); sensitivity analysis shows this across Fréchet and clearance thresholds (Appendix G.2, Fig. 12).
Reasoning-focused models show notable improvements on geometric union tasks, demonstrating that extended inference computation helps with complex geometric operations. However, gains on planning tasks remain modest, suggesting that the bottleneck lies in spatial representation rather than purely computational depth.
Extended capabilities. We also evaluated tool augmentation and vision-language input (Appendix F). Python Code Interpreter yields strong gains on arithmetic-heavy tasks (plus over 30–40 percentage points on Distance/Angle) but limited benefit on planning tasks, indicating failures stem from spatial reasoning rather than calculation errors. Rendered floorplan images provide selective improvements on synthetic layouts but do not consistently outperform symbolic input, suggesting symbolic representations already provide strong baselines.
Future directions. Two complementary directions follow. Near term: hybridize with external geometric solvers or symbolic planning modules, including set operations (unions/differences), centroid via shoelace, clearance buffering, oriented rectangle search, and A* path planning, to compensate for models’ weaknesses in collision avoidance and clearance reasoning. Longer term: train with explicit spatial constraints and harder distributions (irregular, overlap-heavy layouts), and include constraint-violation exemplars and geometry-aware objectives so models learn to maintain coherence under rotation, clearance, and union operations. Beyond these directions, multi-step interaction (e.g., asking an agent to verify or revise its solution using a rendered floorplan) may further improve performance.
FloorplanQA provides a standalone benchmark for measuring progress on spatial reasoning in layout-based tasks, with applications ranging from architectural design to indoor planning.
@@ Acknowledgments
The work is supported by funding from King Abdullah University of Science and Technology (KAUST)—Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program.
@@ Appendix A Technical Appendix and Supplementary Material for FloorplanQA
This appendix provides comprehensive technical details supporting the main paper.
Section B describes synthetic layout generation and Section C covers HSSD extraction procedures. Section D provides detailed statistics on room complexity, object counts, and overlap distributions. Section E presents the complete evaluation breakdown by model, room type, and question category, including truncation analysis. Section F presents extended experiments with tool-augmented reasoning (Python Code Interpreter) and vision-language models using rendered floorplans. Section G provides sensitivity analyses for evaluation tolerances. Section H presents ablation studies examining input format, object reference substitution, and semantic perturbation. Section I contains qualitative failure analysis with annotated examples identifying common error patterns across question types. Section J includes full prompt templates for data generation and question asking. Section K provides supplementary accuracy tables.
@@ B.1 Room Shape Generation
Room Type Shape Distribution (Rect / L-shaped / Open) Size Categories (in m2) (Small / Medium / Large)
@@ Appendix C Details on HSSD layouts selection
During 2D projection, we correct mislabeled windows and doors, snap nearly collinear edges, and resolve self-intersections. To smooth rectilinear artifacts and remove redundant vertices in box-like footprints, we apply Douglas-Peucker polygon simplification with tolerance $\epsilon=0.01$ m. This preserves the polygon’s overall shape while reducing token overhead. All ground-truth answers (centroids, areas, distances) are computed from this cleaned geometry.
@@ Appendix D Room Statistics
Table 4 reports summary statistics for the layouts. Overlaps counts pairs of objects whose bounding boxes intersect. We report two versions: excluding rugs (rugs are ignored when counting overlaps) and including rugs (all overlaps counted). Object density is computed as the total number of objects divided by room area (objects/m²). Vertices per layout includes all polygon vertices from room objects.
Kitchens are typically the smallest spaces, with relatively few objects and overlaps, but a high density due to their compact geometry. Living rooms are the largest, with slightly more objects overall but lower density, reflecting their open layout. Bedrooms fall in between, with similar object counts to kitchens but more frequent overlaps.
@@
Avg. # of Overlaps (excl. rugs) 0.52 1.82 1.52 4.39
Avg. # of Overlaps (incl. rugs) 0.52 6.25 6.2 6.65
@@ E.1 Detailed Accuracy by full dataset
Tables 5 and 6 report the full per-model accuracy matrix for the base and reasoning model groups, respectively. Each table covers nine question categories across four room types, resulting in a total of 36 rows. The Shortest path task is assessed using two complementary metrics: validity and Fréchet Distance. Together, these question types span a range of challenges, covering both reasoning-heavy and functionally grounded tasks. The base group includes eight general-purpose models, while the reasoning group includes seven models with explicit reasoning capabilities. For every model, question type accuracy is computed over a fixed set of 600 synthetic layouts (Kitchens, Living Rooms, Bedrooms) and 200 HSSD layouts, ensuring that results are directly comparable across models.
@@
Question Room [Uncaptioned image] Claude
Sonnet-4
[Uncaptioned image] GPT-4.1
[Uncaptioned image] Kimi-K2
Instruct
[Uncaptioned image] Qwen3
Coder-480B
[Uncaptioned image] Qwen3
235B
[Uncaptioned image] GPT-4.1
mini
[Uncaptioned image] Qwen3
30B
[Uncaptioned image] Devstral
Small
@@
E.2 Token-Limit Analysis and Valid-Only Accuracy
First, Tables 12 and 14 quantify the fraction of responses that were terminated due to the TOKEN LIMIT stop reason. This failure mode is particularly relevant for reasoning-oriented models, which often generate longer chain-of-thought outputs.
Second, Tables 13 and 15 report accuracy over completed answers only, excluding truncated outputs (e.g., token-limit terminations). This metric isolates models’ reasoning performance on successfully produced, well-formed responses.